OpenStudy (anonymous):

Compare the fraction. 1/6 (?) 5/6 A. < B. = C. >

OpenStudy (anonymous):

Compare the fraction. 4/7 (?) 1/6 A. > B. = C. < D. cannot be compared

OpenStudy (anonymous):

to compare two fractions for less than or greater than their denominator must have same values hint: take lcm of denominator

OpenStudy (anonymous):

1.) A 2.) D

hero (hero):

Do you know how to multiply Esmy?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

Esmy what is that? LOL

hero (hero):

Well, you can cross multiply to figure out which is greater

hero (hero):

For example:

OpenStudy (anonymous):

the answer should be A ie 24/42 > 7/42

OpenStudy (anonymous):

A for number one because I know 1 is a

hero (hero):

pulkit21384, you're not helping by giving answers

OpenStudy (anonymous):

@Hero check my previous comment i tried to explain her how it works

hero (hero):

Don't worry, pulknit, I'll explain it

OpenStudy (anonymous):

@Esmy i misread your options option C is the right one

hero (hero):

Stop giving answers or I'm going to report you

hero (hero):

\(\frac{1}{6} ? \frac{5}{6}\) \((1)(6) ? (6)(5)\) 6 < 30 Therefore \(\frac{1}{6} < \frac{5}{6}\)